OpenStudy (anonymous):

george and carmen went on a bicycle trip. they took a bus to their starting point, and then biked the rest. they traveled 200 kilometers in total, and they biked 40 kilometers more than they were bussed. find how many kilometers they traveled by bike

OpenStudy (anonymous):

You need to set up two equations. "They traveled 200 kilometers in total": x + y = 200 "biked 40 kilometers more than they were bussed": 40 + y = x Now solve by substitution: 40 + y + y = 200 2y = 160 y = 80 Sub y = 80 to x + y = 200 x + 80 = 200 x = 120 Thus George and Carmen biked for 120 km and bussed for 80 km.